the production function is Q = 16K0.75 L1.25. Given that price of capital is EC$32 and the MRTS is K = 2L a) what is the price of labour b) if the desired output is 8,500 units, how many units of capital and labour will be used the subject is managerial economics
Answer #1

0= 16 K075625 MPL = delar MPL = (16 (105) Kor 123 IMPL 20 Kars 20.257 MPK = de lok NPK (16) (0.75) K 0.1-1, 1.25 I MYK - 12 Kcha 1640.75, 1.25 put = 850. 25 >8560 = 16 kuns put K = 2L - 850= 16 (22) 0.75 11-25 ... - 8500 = (16) (2) 0.15 10.15 21:25 8